OpenStudy (dumbsearch2):

Together, Tim and Alan can unload a truck 6 in minutes. When Tim works alone, it takes him 9 minutes longer than Alan to unload the truck. How long does it take each person to unload the truck alone? So, I simplified the denominator, So I had this:T(T+9)/2T+9=6. Then, I multiplied the denominator on both sides. T(T+9)=6(2T+9). I distributed.T(T+9)=12T+54 Then...Now here is the part I am not sure about...I distribute the T: T^2+9T=12T+54. I moved the 9T to the right side: T^2=12T-9T+54. So now I have this: T^2=3T+54 I am not sure what to do next. How would I go about it? thanks!!!!

OpenStudy (dumbsearch2):

So I found 9 for Tim after solving the quadratic. But I am sort of confused about Alan. If I understand correctly, T=9+A. But if this is the case, then A is 0. Is this even possible? Thanks again for your help. :)

OpenStudy (mertsj):

\[\frac{1}{t}+\frac{1}{t+9}=\frac{1}{6}\] \[6t+54+6t=t^2+9t\] \[t^2-3t-54=0\] \[(t-9)(t+6)=0\] t=9=Alan's time t+9=18 = Tim's time